All discount rates are 6% per period unless otherwise stated. All cash flows are end-of-period. Answers are in parens, from Excel, so they may differ slightly from table answers due to rounding.

  • $1000 per period is deposited (at the end of the period) in an account earning 8%/period, for 10 periods (years). Additionally, starting in Y2, an additional $200 is deposited, increasing each year. (So CF’s are $1000, 1200, 1400, 1600...). What’s the final value(end of 10th period)? ($25,702.97)

  • $2000 per year is deposited in an account earning 6%/year. Starting in Y5, an extra $100, increasing annually, is deposited. (Y1,2,3,4=2K; Y5=2100, Y6=2200...). What’s the value at end of year 11? ($33,105.73)
Answer #1

a. Future worth of cash flows = 1000 * (F/A,8%,10) + 200 * (F/G,8%,10)

= 1000 * 14.486562 + 200 * 56.082031

= 25702.97

b. Future worth of cash flows = 2000 * (F/A,6%,11) + 100 * (F/G,6%,8)

= 2000 * 14.971643 + 100 * 31.624465

= 33105.73